OpenStudy (anonymous):

Which line shows the first error in the solution? 8x-6=4x-2 (1) 8x-6=4x-4x-2 (2) 4x-6=2 (3) 4x=6+6=2=6 (4) 4x=8 (5) 4x/4=8/4 (6) x=2 (7)

OpenStudy (zpupster):

8x-6=4x-2 ok 8x-6=4x-4x-2 subtract 4x both sides- shown only one 4x-6=2 answer is ok 4x=6+6=2=6 lot wrong here